If a train runs at 40 kmph, it reaches its destination late by 11 minutes but if it runs at 50 kmph, it is late by 5 minutes only. The correct time for the train to complete its journey is:
Answer by richwmiller(17219)

40(t+11)=d,
50(t+5)=d
d = 1200, t = 19
at 40 kmh it should go for 30 minutes 40*30=1200
at 50 kmh it should go for 24 minutes 50*24=1200